I'm really lost on how to do these.
Depreciation,solve the literal equation for the indicated variable.
D = c- S
______ (for s)
N


Answer by stanbon(48535)

Solve D = (c- s)/N for s

Cross-multiply to get:
DN=c-s
Then s=c-DN
Cheers,
Stan H.